OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 5 A.x + (y + 5) = 6 B.x + y = y + 5 C.x = (x – 6) + 5 D.(y + 5) + y = 6

OpenStudy (anonymous):

I believe its is d

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

[1] x+y=6 [2] x=y+5 substitution = moving one equation into the other. since [2] x=y+5 put [2] into [1] x+y=6 replacing x with y+5 and that is why

OpenStudy (anonymous):

Thank you for explaining it I had a big problem with this. I had to go to my teacher and get help.